1-Question 15 3.03 pts Write the function in terms of its cofunction. Assume that any angle in which an unknown appears is an a

2-Question 18 3.03 pts Find the quotient and write in rectangular form. First convert the numerator and denominator to trigonom
Answer #1

SOLUTION;

15)

tan(\theta+60)

=tan(\theta+60+30-30)

=tan(\theta+90-30)

=tan(90+\theta-30)

=tan(90 -(30-\theta))

=cot(30-\theta)

Answer

=cot(30-\theta)